A 830 kg car travels around a curve of 180 m radius at 80 km/h. The acceleration of gravity is 9.81 m/s^2.        What should be the banking angle of the curve so that the force of the pavement on the tires of the car is in the normal direction? Answer in units of degrees.
Answer #1

convert 80 km/h in SI sytem:
80(1000)/3600 = 22.22 m/s

Draw Free-body diag of car on incline =>show Fc (centripetal force), weight = mg & Normal force.

? = angle of bank of curve/s
tan ? = Fc/mg = mV
Answer #2

tan(?)=(v^2)/rg

23 degree
Answer #3

theta = arctan ( v^2 / rg)

= arctan ( (80*5/18)^2 / (180*9.81))

= 15.624 degrees